In a class of 330 students, there are 60 more girls than boys. How many girls are there in the class?
Answer by DrBeeee(684)

Let g = the number of girls in the class
Let b = the number of boys in the class
Given:
(1) g + b = 330
and
(2) g = b + 60
Put g of (2) into (1) and get
(3) b + 60 + b = 330 or
(4) 2b + 60 = 330 or
(5) 2b = 330 - 60 or
(6) 2b = 270
Now divide each side of (6) by 2 to get
(7) 2b/2 = 270/2 or
(8) b = 135
Then from (2) we get
(9) g = 135 + 60 or
(10) g = 195
Let's check these values with (1).
Is (195 + 135 = 330)?
Is (330 = 330)? Yes
Answer: There are 195 girls in the class.